Pre-Algebra Chapter 3 Resource Book Lesson 3.3 Practice B (Don't know if this will help)
Tell whether the given value of the variable is a solution of the equation.
question 2:
4x + 13 = -9 -3(x+9); x = -7

Answer by bucky(2189)

Given:
.
4x + 13 = -9 -3(x + 9)
.
To see if x = -7 is a solution of this equation, all you need to do is to substitute -7
for x in the equation. Then if the left side equals the right side, -7 is a solution.
However, if the two sides are unequal, then it is not a solution.
.
So let's begin with the given equation:
.
4x + 13 = -9 - 3(x + 9)
.
substitute -7 for x and the equation becomes:
.
4*(-7) + 13 = -9 -3(-7 + 9)
.
On the left side multiply 4 times -7 to get -28. The equation then becomes:
.
-28 + 13 = -9 - 3(-7 + 9)
.
On the left side combine the -28 and the +13 to get -15. This reduces the equation to:
.
-15 = -9 - 3(-7 + 9)
.
On the right side combine the two numbers in the parentheses. -7 plus 9 is +2 and this
changes the right side to:
.
-15 = -9 - 3(2)
.
Multiply the -3 times +2 and it becomes -6, making the equation:
.
-15 = -9 - 6
.
Combine the two numbers on the right side and you have:
.
-15 = -15
.
When x equals -7 the two sides of the equation are each -15 and therefore are equal to
each other. This means that x = -7 is a solution to the equation because it makes both
sides of the equation equal to each other.
.
Hope this helps you to see how to do this problem.
.